# Qingyu Shi1  Lu Qi2  Jianzong Wu1   Jinbin Bai3

### v2

@@ Abstract
Customized image generation is crucial for delivering personalized content based on user-provided image prompts, aligning large-scale text-to-image diffusion models with individual needs. However, existing models often overlook the relationships between customized objects in generated images. Instead, this work addresses that gap by focusing on relation-aware customized image generation, which aims to preserve the identities from image prompts while maintaining the predicate relations described in text prompts. Specifically, we introduce RelationBooth, a framework that disentangles identity and relation learning through a well-curated dataset. Our training data consists of relation-specific images, independent object images containing identity information, and text prompts to guide relation generation. Then, we propose two key modules to tackle the two main challenges—generating accurate and natural relations, especially when significant pose adjustments are required, and avoiding object confusion in cases of overlap. First, we introduce a keypoint matching loss that effectively guides the model in adjusting object poses closely tied to their relationships. Second, we incorporate local features from the image prompts to better distinguish between objects, preventing confusion in overlapping cases. Extensive results on three benchmarks demonstrate the superiority of RelationBooth in generating precise relations while preserving object identities across a diverse set of objects and relations. The source code and trained models will be made available to the public.
Figure 1: Relation-aware image customization, the generated image should strictly keep the predicate relation and preserve each identity among those identities provided by the text prompt.
@@ 1 Introduction
Driven by large-scale text-to-image diffusion models (Rombach et al., 2022; Saharia et al., 2022; Podell et al., 2023; Pernias et al., 2023), customized image generation has recently made significant strides (Ruiz et al., 2023; Kumari et al., 2023; Wei et al., 2023; Zhang et al., 2024; Wang et al., 2024; Kong et al., 2024; Bai et al., 2024). This task focuses on generating images that preserve the identity of objects from user-provided inputs, enabling the creation of personalized and meaningful content. It has shown value in numerous applications, including personalized artwork, branding, virtual fashion try-ons, social media content creation, augmented reality experiences, and more.
@@
Despite the success of many methods for customizing single or multiple objects (Liu et al., 2023; Ye et al., 2023; Li et al., 2023; Patel et al., 2024; Wang et al., 2024; Gu et al., 2024; Lin et al., 2024; Pang et al., 2024), they often overlook the relationships between objects and the corresponding text prompts. For instance, when two user-provided objects are paired with a text prompt specifying a particular relationship, the generated output should not only preserve their identities but also accurately reflect the intended relationship, such as a ‘hug’. This introduces new challenges in what we refer to as relation-aware customized image generation, which focuses on preserving multiple identities while adhering to relationship prompts.
@@
An intuitive solution to this issue is to adapt existing tuning-based or training-based methods for customizing relationships (Huang et al., 2023; mengmeng Ge et al., 2024; Materzynska et al., 2023) among objects. However, both approaches face challenges. Tuning-based methods struggle to preserve multiple identities, as they invert objects into specific tokens, while training-based schemes often fail to balance image and text prompts, frequently overlooking key textual elements and hindering the generation of relationships between objects. As illustrated in Fig. 1, previous methods fail to capture the actions described in the text prompt and lose identity preservation.
@@
We attribute this failure to two key factors: a lack of relevant data and an ineffective model design. Unlike data augmentation techniques such as flipping or rotation, commonly used to create paired training data in object customization methods (Yang et al., 2023; Chen et al., 2024b), our approach requires a triplet of images: two image prompts and one target image. The image prompts should contain similar objects but exhibit distinct actions compared to the target image. To collect these triplets, we propose a data engine to curate our fine-tuning set. We leverage an advanced text-to-image generation model (Betker et al., 2023) to generate triplets where the same object pair is shared across the images. Through text prompt guidance, the object in the image prompt provides strong identity information while maintaining distinct actions in the target image, enabling the model to focus on relation learning during fine-tuning.
@@
For the model design, we propose RelationBooth, which applies the Low-Rank Adaptation (LoRA) (Hu et al., 2022) strategy to the text cross-attention layers of existing diffusion models to process user-provided text prompts. In RelationBooth, two key modules are introduced during training to enhance the customization of relationships and identity preservation. First, we introduce a keypoint matching loss (KML) as additional supervision to explicitly encourage the model to adjust object poses, since relationships between objects are closely tied to their poses. Importantly, the KML operates on the latent representation rather than the original image space, aligning with the default diffusion loss. Second, we inject local tokens for multiple objects to improve the distinctiveness of highly overlapping objects. Specifically, we employ the self-distillation method from CLIPSelf (Wu et al., 2024b) to enhance region-language alignment in CLIP’s dense features. Through partitioning and pooling, the well-aligned local tokens help mitigate appearance confusion between objects.
@@
To more comprehensively evaluate relation-aware customized image generation, we constructed our RelationBench based on three established benchmarks (Ruiz et al., 2023; Kumari et al., 2023). Our methods demonstrate strong performance compared to existing approaches, achieving significant improvements in visual quality and quantitative results.
@@
  • We explore a novel task called relation-aware customized image generation, which aims to preserve multiple identities from image prompts while adhering to the relationships specified in text prompts. This task can enhance various user-driven applications by enabling more control and customization.
  • We introduce a data engine that uses an advanced text-to-image generation model to generate triplet images where the same object pair is present with distinct actions. This well-curated dataset enables the model to focus on learning object relations during fine-tuning, leveraging the robust identity information provided by the image prompts.
  • Our proposed RelationBooth method has two key modules, including keypoint matching loss and object-related local tokens, to enhance the customization of relationships and identity preservation. The extensive experiments on three benchmarks demonstrate the effectiveness of our method.
@@ 2 Related Work
Diffusion-Based Text-to-Image Generation. This direction aims to produce high-quality images based on user-provided textual prompts. Diffusion-based generation models (Rombach et al., 2022; Saharia et al., 2022) encode these prompts through the text model (Radford et al., 2021; Chung et al., 2022) and inject the text embedding into the U-Net’s cross-attention layers. Some methods (Podell et al., 2023; Pernias et al., 2023; Ren et al., 2024; Wu et al., 2024a) upscale diffusion models and incorporate additional conditions as priors to generate high-resolution images. Meanwhile, several methods (Betker et al., 2023; Lian et al., 2023) introduce stronger text encoders or large language models (LLMs) to enhance the textual comprehension capabilities of diffusion models. Recent works (Peebles & Xie, 2023; Chen et al., 2024a) have replaced the U-Net denoiser with a Transformer-based denoiser. However, these models primarily focus on textual conditions and struggle to handle other forms of input, such as user-provided images. In contrast, our work facilitates customized generation while following predicate relation conditions given by text inputs.
@@
Diffusion-Based Customized Generation. Customized image generation aims to produce diverse images based on user-provided concepts. Tuning-based methods accomplish this by fine-tuning specific parameters of diffusion models, thereby incorporating new concepts into text-to-image diffusion models like Stable Diffusion. Some of these methods (Gal et al., 2023; Voynov et al., 2023; Liu et al., 2023) employ text embeddings to represent customized concepts, which are then injected into prompts during inference. On the other hand, Dreambooth (Ruiz et al., 2023) fine-tunes the entire U-Net and introduces a prior preservation loss to mitigate language drift. Additionally, several approaches (Kumari et al., 2023; Gu et al., 2024; Lin et al., 2024; Pang et al., 2024) combine text embeddings and diffusion models for image synthesis. They fine-tune both the text embedding and the U-Net, yielding impressive results. Considering the substantial cost of fine-tuning for commercial uses, training-based methods (Li et al., 2023; Wei et al., 2023; Chen et al., 2024b; Patel et al., 2024; Zhang et al., 2024) have been proposed. This approach primarily utilizes an ID encoder to extract the concept’s identity and inject it into the U-Net via cross-attention layers. For instance, MS-Diffusion (Wang et al., 2024) integrates grounding tokens with a feature resampler to maintain detail fidelity and employs layout guidance to explicitly place the concepts. Despite their efficiency during inference, training-based methods often struggle with identity preservation and textual condition following ability. A typical problem is the relation expressions in prompts are often overlooked. To address this issue, we present RealtionBooth, to tackle relation-aware customized generation.
@@
Relation-Aware Text-to-Image Generation. Inspired by tuning-based customization methods (Gal et al., 2023), some recent works (Huang et al., 2023; Wei et al., 2024; mengmeng Ge et al., 2024) represent a “neglected word” by learnable parameters. These methods fine-tune part of the parameters on content co-existing images. For instance, Reversion (Huang et al., 2023) fine-tunes the text embedding on a set of co-existing related images and introduces a relation-steering contrastive loss. Reversion gains a better alignment between relational words and generated images by injecting new text embedding into prompts. While effective with prepositions and adjectives, Reversion struggles with predicate relation which involves significant overlaps, and cannot customize user-provided concepts. Moreover, this method still struggles to generate vivid relationships while maintaining the fidelity of multiple customized concepts, likely due to the inherent disregard for text embeddings in customization methods. Our work bridges the gap in the predicate relation-following capability of training-based customization methods. This enables efficient inference that naturally and accurately generates predicate relations from the textual prompt.
3 Relation-aware customized obejct generation
@@
Figure 2: The overview of RelationBooth. RelationBooth encodes the image prompts by dual encoder architecture. After getting the U-Net output $\hat{\epsilon}$, we predicate $\hat{z}_{0}$ and calculate the keypoint matching loss. The part in the dotted box is only for training.
@@ 3.1 Problem definition
Different from convention customization tasks, we explore a new setting that focuses on image generation by both image prompts $c_{i}\in\mathcal{R}^{N\times H\times W}$ and text prompts $c_{t}$. The $N$, $H$, and $W$ are the number, height, and width of image prompts. We call this setting RelationBooth due to the requirements that the generated image $\hat{x}$ should strictly preserve each identity and keep the predicate relation among those identities provided by $c_{i}$ and $c_{t}$.
@@
where $\Phi$ is the network parametered by $\theta$. For brevity clarification, we set $N=2$ due to the basic triplet element defined in the RelationBooth.
3.2 Discussion and Motivation of RelationBooth
Although previous works (Kumari et al., 2023; Gu et al., 2024; Zhang et al., 2024) have attempted to customize multiple objects under text control, we are the first to address the generation of relation-aware customized images. Notably, the state-of-the-art multi-object customization model, MS-Diffusion (Wang et al., 2024), cannot handle this task effectively. Specifically, given the image prompt $c_{i}$ and text prompt $c_{t}$. MS-Diffusion utilizes the CLIP (Radford et al., 2021) to extract the image and text tokens from $c_{i}$ and $c_{t}$ respectively. The image and text tokens are injected into the U-Net $\bm{\epsilon_{\theta}}$ through parallel cross-attention layers. The U-Net is optimized on the latent representation of image $z$:
$\mathcal{L}_{MS}=\mathbb{E}_{z_{t},t,\epsilon,c}[\|\epsilon_{t}-\bm{\epsilon_{\theta}}(z_{t},t,c_{i},c_{t})\|_{2}^{2}]$ (2)
where $\epsilon_{t}\sim\mathcal{N}(0,1)$, we omit the CLIP encoder for brevity. However, despite MS-Diffusion can customize objects under text control, it’s struggled for MS-Diffusion to generate relations accurately. Furthermore, MS-Diffusion might cause object confusion when heavy overlapping exists, leading to failure relation generation.
We revise the relation generation task. Existing relation inversion methods (Huang et al., 2023; mengmeng Ge et al., 2024) aim to invert a relation to a text embedding $R^{*}$ in the pre-trained text-to-image diffusion models. Given a set of relation-specific images $\{x_{k}\}_{k=1}^{n}$, they employ denoising loss $\mathcal{L}_{RI}$ to fine-tune $R^{*}$ for alignment with specific relation:
@@ 3.2 Discussion and Motivation of RelationBooth
Directly implementing relation inversion methods on MS-Diffusion leads to suboptimal results. We fine-tune MS-Diffusion on a set of relation-specific images, and experiment results show the model falls short in generating relation in Fig. 4. We think the image prompt $c_{i}$ that is not considered in relation inversion methods affects the generation process, leading to an overlook of $R^{*}$. In the following subsections, we will introduce the data collection and proposed method, considering lacking both of them in the proposed tasks.
@@ 3.3 Data Collection
Considering the scarcity of available datasets that can be used, we first design a data engine for collecting high-quality tuning data. We denote the relation-specific image as $x$. While $x$ contains relation information, it also includes the object’s identities. Directly fine-tuning on a set of $x$ leads to unsatisfied relation generation. To disentangle the relation and identity in $x$. The ideal tuning data should be in the form of $\mathcal{D}=(x,c_{i},c_{t})$. Where $c_{i}$ represents the image prompts that contain the object in $x$, $c_{t}$ represents the text prompt that guides the relation generation. Unfortunately, directly cropping the objects from $x$ to obtain $c_{i}$ results in a copy-and-paste effect, even when using data augmentation techniques like flipping and rotation. Therefore, we introduce our data engine below.
@@
Data engine. Fig. 3 illustrates our data generation process. Inspired by the capability of DALL-E 3 and its multi-turn dialog capability. We use DALL-E 3 to generate $\mathcal{D}=(x,c_{i},c_{t})$. Where $x$ and $c_{i}$ share the same object identity by leveraging the prompt “The photo of the same.” We observe that prompting in this manner allows DALL-E 3 to remember and preserve identity in common categories such as “tiger” and “brown bear” which are enough for relation learning. Next, we employ X-Pose (Yang et al., 2024), SAM (Kirillov et al., 2023), and LLaVA (Liu et al., 2024) to annotate $x$, $c_{i}$ with keypoints, masks, and captions.
Figure 3: The overview of the data generation process. DALLE-3 can generate similar objects by using ’The photo of the same’ in the text prompt on common animal categories.
Figure 4: Comparison of different fine-tuning data. Simply implementing relation inversion methods on MS-Diffusion leads to failure relation generation.
3.4 RelationBooth
3.4.1 RelationLoRA Fine-tuning
To disentangle the relation and identity information in relation-specific images $x$, we utilize off-the-shelf identity inject methods from state-of-the-art customized model MS-Diffusion (Wang et al., 2024). Specifically, we utilize parallel cross-attention layers to process the $c_{i}$ and $c_{t}$, the identity information in $c_{i}$ facilitates the model to focus on relation information in $x$ and $c_{t}$ during fine-tuning. We formulate the parallel cross-attention layers:
$h=\gamma\cdot\operatorname{Softmax}(\frac{QK_{i}}{\sqrt{d}})V_{i}+\operatorname{Softmax}(\frac{QK_{t}}{\sqrt{d}})V_{t}$ (4)

### v3

# DreamRelation: Bridging Customization and Relation Generation

@@ Abstract
Customized image generation is essential for creating personalized content based on user prompts, allowing large-scale text-to-image diffusion models to more effectively meet individual needs. However, existing models often neglect the relationships between customized objects in generated images. In contrast, this work addresses this gap by focusing on relation-aware customized image generation, which seeks to preserve the identities from image prompts while maintaining the relationship specified in text prompts. Specifically, we introduce DreamRelation, a framework that disentangles identity and relation learning using a carefully curated dataset. Our training data consists of relation-specific images, independent object images containing identity information, and text prompts to guide relation generation. Then, we propose two key modules to tackle the two main challenges—generating accurate and natural relationships, especially when significant pose adjustments are required, and avoiding object confusion in cases of overlap. First, we introduce a keypoint matching loss that effectively guides the model in adjusting object poses closely tied to their relationships. Second, we incorporate local features of the image prompts to better distinguish between objects, preventing confusion in overlapping cases. Extensive results on our proposed benchmarks demonstrate the superiority of DreamRelation in generating precise relations while preserving object identities across a diverse set of objects and relationships. The source code and trained models will be made available to the public.
11footnotetext: * Corresponding authors: Lu Qi, Xiangtai Li.
Figure 1: In our Relation-Aware Image Customization task, the generated images must accurately preserve the relationships between objects and maintain their identity. We highlight the limitations of previous approaches using three color codes: red indicates failure to capture relationships, blue marks missing objects, and orange represents object confusion. Each image is annotated to reflect its specific issue. Our results, highlighted by green boxes, demonstrate the advantages of our proposed method.
@@ 1 Introduction
Driven by large-scale text-to-image diffusion models [30, 32, 27, 26], customized image generation has recently made significant strides [31, 15, 36, 43, 35, 14, 1]. This task focuses on generating images that preserve the identity of objects from user-provided inputs, enabling the creation of personalized and meaningful content. It has shown value in numerous applications, including personalized artwork, branding, virtual fashion try-ons, social media content creation, augmented reality experiences, and more.
@@
Despite the success of many methods for customizing single or multiple objects [20, 42, 16, 24, 35, 8, 18, 23], they often overlook the relationships between objects and the corresponding text prompts. For instance, when two user-provided objects are paired with a text prompt specifying a particular relationship, the generated output should not only preserve their identities but also accurately reflect the intended relationship, such as a ‘hug’. This introduces new challenges in what we refer to as relation-aware customized image generation, which focuses on preserving multiple identities while adhering to relationship prompts.
@@
An intuitive solution to this issue is to adapt existing relation-aware generation methods [11, 22, 21] to tuning-based or training-based customized methods for customizing objects while maintaining the relation. However, both approaches face challenges. Tuning-based methods struggle to preserve multiple identities, as they invert objects into specific tokens. While training-based schemes often fail to balance image and text prompts, frequently overlooking key textual elements and hindering the generation of relationships between objects. As illustrated in Fig. 1, previous methods fail to capture the relationships described in the text prompt and lose identity preservation.
@@
We attribute this failure to two key factors: a lack of relevant data and an ineffective model design. Unlike data augmentation techniques such as flipping or rotation, commonly used to create paired training data in object customization methods [40, 5], our approach requires a triplet of images: two image prompts and one target image. The image prompts should contain similar objects but exhibit distinct poses compared to the target image. To collect these triplets, we propose a data engine to curate our fine-tuning set. We leverage an advanced text-to-image generation model [2] to generate triplets where the same object pair is shared across the images. Through text prompt guidance, the image prompt provides strong identity information. This enables the decoupling of the relationship in the target image, which enhances the relation learning.
@@
For the model design, we propose DreamRelation, which applies the Low-Rank Adaptation (LoRA) [9] strategy to the text cross-attention layers of existing diffusion models to process user-provided text prompts. In DreamRelation, two key modules are introduced during training to enhance the relation generation in customized generation. First, we introduce a keypoint matching loss (KML) as additional supervision to explicitly encourage the model to manipulate object poses, since relationships between objects are closely tied to their poses. Importantly, the KML operates on the latent representation rather than the original image space, aligning with the default diffusion loss. Second, since relation-aware customization requires local features from image prompts, such as the “hands” features for generating “shaking hands”—which are not captured by CLIP’s coarse image-level features, we introduce dense features from CLIP. Through partitioning and pooling, we obtain local tokens that contain detail and local information. To further enhance the compatibility between dense features and image-level features, we employ a self-distillation method [39] to improve their alignment.
@@
To more comprehensively evaluate relation-aware customized image generation, we constructed our RelationBench based on three established benchmarks [31, 15]. Our methods demonstrate strong performance compared to existing approaches, achieving significant improvements in visual quality and quantitative results.
@@
  • We explore a novel task called relation-aware customized generation, which aims to preserve multiple identities from image prompts while adhering to the relationships specified in text prompts. This task can enhance various user-driven applications by enabling more control.
  • We introduce a data engine that uses the recent text-to-image generation model to generate triplet images where the same object is present with distinct poses. This well-curated dataset enables the model to focus on relation learning during fine-tuning, leveraging the identity information provided by the image prompts.
  • Our proposed DreamRelation has two key modules, including keypoint matching loss and local token injection, to enhance the relation generation in customized generation. The extensive experiments on three benchmarks demonstrate the effectiveness of our method.
@@ 2 Related Work
Diffusion-Based Text-to-Image Generation. Diffusion-based text-to-image models [30, 32] generate high-quality images based on user-provided text prompts. These works encode the text prompt through the text encoder [28, 6] and inject the text embedding into the U-Net’s cross-attention layers. Some methods [27, 26, 29, 38] upscale diffusion models and incorporate additional conditions as priors to generate high-resolution images. Meanwhile, several methods [2, 17] introduce stronger text encoders or large language models (LLMs) to enhance the text comprehension capabilities of diffusion models. Recent works [25, 4] have replaced the U-Net denoiser with the Transformers [33]. However, these models primarily focus on text conditions and struggle to handle other forms of guidance, such as user-provided images. In contrast, our work facilitates customized generation while following relation conditions given by text inputs.
@@
Diffusion-Based Customized Generation. Customized image generation aims to produce new images based on user-provided objects. Tuning-based methods accomplish this by fine-tuning specific parameters of diffusion models to learn new objects. Some of these methods [7, 34, 20] employ text embeddings to represent the object identity. On the other hand, DreamBooth [31] fine-tunes the entire U-Net and introduces a prior preservation loss to mitigate language drift. Additionally, several approaches [15, 8, 18, 23] integrate text embeddings and U-Net parameters for image customization. These methods fine-tune both components during object learning, achieving impressive results. Considering the substantial cost of fine-tuning for commercial uses, training-based methods [16, 36, 5, 24, 43] have been proposed. This approach primarily utilizes an encoder to extract the object’s identity and inject it into the U-Net. For instance, MS-Diffusion [35] integrates grounding tokens with a Resampler to maintain detailed identity and employs layout guidance to explicitly locate the objects. Despite their efficiency during inference, training-based methods often face challenges in balancing identity preservation and text control. A common issue is that the relationships described in text prompts are frequently overlooked. To address this limitation, we introduce DreamRelation, a framework designed to bridge the customization and relation generation.
@@
Relation-Aware Text-to-Image Generation. Inspired by tuning-based customization methods [7], some recent works [11, 37, 22] represent a “neglected word” by learnable parameters. These methods fine-tune part of the parameters on the content co-existing images. For instance, Reversion [11] fine-tunes the text embedding on a set of relation co-existing images and introduces a relation-steering contrastive loss. Reversion gains a better alignment between relational words and generated images by injecting new text embedding into prompts. While effective with prepositions and adjectives, Reversion struggles with relationships that involve significant overlaps, and cannot customize user-provided objects. Moreover, this method still struggles to generate vivid relationships while maintaining the fidelity of multiple customized objects, likely due to the inherent disregard for text embeddings in customization methods. Our work bridges the gap in the predicate relation-following capability of training-based customization methods. This enables efficient inference that naturally and accurately generates relationships in the text prompts.
3 Bridging Customization and Relation Generation
@@
Figure 2: The overview of DreamRelation. DreamRelation utilizes the off-the-shelf identity extractor to decouple the relation and identity information in relation-specific images. After getting the U-Net output $\hat{\epsilon}$, we predict $\hat{z}_{0}$ and calculate the keypoint matching loss. The part in the dotted box is only for training.
@@ 3.1 Problem definition
Different from conventional customization tasks, we explore a new setting that focuses on image generation by both image prompts $c_{i}\in\mathcal{R}^{N\times 3\times H\times W}$ and text prompts $c_{t}$. We call this setting relation-aware customized image generation due to the requirements that the generated image $\hat{x}$ should strictly preserve each identity and keep the relationship among those identities provided by $c_{i}$ and $c_{t}$.
@@
where $\Phi_{\theta}$ is the network parameterized by $\theta$. For brevity clarification, we set $N=2$ due to the basic triplet element defined in the DreamRelation.
3.2 Discussion and Motivation
While previous works [15, 8, 43] have explored the customization of multiple objects under text control, they lack the ability to effectively manage the relationships between these objects. Specifically, the state-of-the-art multi-object customization model [35] utilizes bounding boxes to enhance multi-object generation. Given the image prompt $c_{i}$ and the text prompt $c_{t}$, it leverages CLIP [28] and the Resampler [35] to extract text tokens $tok_{\text{text}}$ and image tokens $tok_{\text{image}}$, respectively. These tokens are then injected into the U-Net $\bm{\epsilon_{\theta}}$ through parallel cross-attention layers:
$h=\text{Softmax}(\frac{QK_{i}}{\sqrt{d}}+\mathcal{M})V_{i}+\text{Softmax}(\frac{QK_{t}}{\sqrt{d}})V_{t}$ (2)
where $Q$ represents the query, while $K_{i},V_{i}$ and $K_{t},V_{t}$ denote the keys and values obtained from $tok_{\text{image}}$ and $tok_{\text{text}}$, respectively. The mask $\mathcal{M}$, generated from the bounding boxes, assists the model in localizing objects during multi-object customization. However, it encounters difficulties when bounding boxes overlap, frequently resulting in object confusion, as shown in Fig. 1. Consequently, relation-aware generation in customized models remains a largely unexplored area of research.
On the other hand, we revisit the relation-aware generation task. Existing relation inversion methods [11, 22] focus on inverting a relationship into a text embedding $R^{*}$ within pre-trained text-to-image diffusion models. Given a set of images $\{x_{k}\}_{k=1}^{n}$ that share the same relationship, these methods use a denoising loss $\mathcal{L}_{RI}$ to fine-tune $R^{*}$, ensuring alignment with the specific relationship:
@@ 3.2 Discussion and Motivation
Directly applying relation-aware generation methods to customized models leads to suboptimal performance, as illustrated in Fig. 4. We attribute this to two main reasons. First, the balance between the control of $c_{i}$ and $c_{t}$ is not addressed in relation-aware generation tasks. During inference, the model tends to overlook the target relationship $R^{*}$ due to the influence of $c_{i}$, resulting in incorrect relationship generation. Second, even if the shared relationship exists across $\{x_{k}\}_{k=1}^{n}$, the disentangled object identity information complicates the learning process. Fine-tuning solely on $\{x_{k}\}_{k=1}^{n}$ makes it challenging for the model to effectively capture the relationship. As shown in the second row of Fig. 4, where ReVersion [11] struggles to generate the ”hugging” relationship even in the absence of $c_{i}$. In the following subsections, we will introduce our data collection method and DreamRelation to bridge customization and relation generation.
@@ 3.3 Data Collection
To address the balance between $c_{i}$ and $c_{t}$ while facilitating relation learning from $\{x_{k}\}_{k=1}^{n}$, we first design a data engine for collecting high-quality tuning data. The ideal tuning data should be in the form of $\mathcal{D}=(x_{k},c_{i},c_{t})$. The $c_{i}$ contain the object in $x_{k}$ while $c_{t}$ describes the relationship between objects in $x_{k}$. The identity information in $c_{i}$ helps decouple the relationship within $x_{k}$, enhancing relation learning while maintaining a balanced integration of $c_{i}$ and $c_{t}$. Unfortunately, directly cropping the objects from $x_{k}$ to obtain $c_{i}$ results in a copy-and-paste effect [5], even when using data augmentation techniques like flipping and rotation. Therefore, we introduce our data engine below.
@@
Relation-aware data engine. Fig. 3 illustrates the difference of our data engine. Both of $x_{k}$ and $c_{i}$ are generated by the recent text-to-image generation model, rather than through cropping and augmentation. Specifically, inspired by the performance of DALL-E 3 and its multi-turn dialog capability. We use DALL-E 3 to generate $\mathcal{D}=(x,c_{i},c_{t})$. Where $x_{k}$ and $c_{i}$ share the same object identity by leveraging the prompt “The photo of the same.” We observe that prompting in this manner allows DALL-E 3 to remember and preserve identity in common categories such as ”tiger”, which is enough for relation learning. Next, we employ X-Pose [41], SAM [13], and LLaVA [19] to annotate $x$, $c_{i}$ with keypoints, masks, and captions for samples.
Figure 3: The rigid object cropping data engine, which results in minimal changes to object pose, the cropped image prompts also contain relation information, leading to a copy-and-paste effect, where the text prompt is neglected. Our relation-aware data engine, on the other hand, focuses on relation learning by decoupling identity information from images.
3.4 DreamRelation
3.4.1 Relation Learning
To balance $c_{i}$ and $c_{t}$ while enhancing relation learning, we decouple the identity information and relationships in $x_{k}$. Specifically, we leverage an off-the-shelf ID extractor from the state-of-the-art customized model [35]. The identity information in $c_{i}$ enables the model to focus on the relationship information in $x_{k}$ and $c_{t}$ during relation learning. To achieve this, we employ parallel cross-attention layers to process $c_{i}$ and $c_{t}$, which can be formulated as:
$h=\gamma\cdot\text{Softmax}(\frac{QK_{i}}{\sqrt{d}}+\mathcal{M})V_{i}+\text{Softmax}(\frac{QK_{t}}{\sqrt{d}})V_{t}$ (4)